OpenStudy (loser66):

Suppose the series \(\sum_{n=1}^\infty a_nx^n\) converges for x =-3 and diverges for x =5 a) What can you conclude about the radius of convergence of the series? b) Does \(\sum_{n=1}^\infty a_n(-6)^n \) converge? Does \(\sum_{n=1}^\infty na_n2^n \) converge? What do you think?

OpenStudy (loser66):

@SithsAndGiggles I know you are not online, just in case.

OpenStudy (loser66):

I've just finish my final test. This is the problem I got stuck.

OpenStudy (anonymous):

By the ratio test, the series would converge if \[|x|\lim_{n\to\infty}\left|\frac{a_{n+1}}{a_n}\right|<1\] Since this is true for \(x=-3\), this means we have a minimum radius of convergence of \(3\). Also, given that the series diverges for \(x=5\), we know that the maximum radius must be \(5\).

OpenStudy (loser66):

We are not allowed to take ratio test.:(

OpenStudy (loser66):

However, it is not <1

OpenStudy (anonymous):

The ratio test isn't the important part here. The given series is centered about \(x=0\), right? The fact that it converges for \(x=-3\) is enough to know that the radius of convergence must be at least \(3\).

OpenStudy (loser66):

This is what I tried: at x =-3, the sum is \(\sum_{n\rightarrow \infty} a_n (-3)^n\) converges, that shows \(lim_{n\rightarrow \infty}a_n = 0\)

OpenStudy (loser66):

not, lim of the whole thing \